# Testing the Froggatt-Nielsen Mechanism with Lepton Violation

### v1

@@ A.3 Predicting Lepton Violation in FN
For the predictions, we employ the SMEFT framework. We construct effective operators following the power counting dictated by formal $U(1)_{\rm FN}\,$invariance, as in Eqs. (17)-(19). We assume all Wilson coefficients to be $\mathcal{O}(1)$ , and generate them randomly with the same priors used for the effective Yukawa coefficients. For simplicity, we consider only SMEFT operators that contribute at tree-level to our observables (we checked explicitly with wilson [41] that running effects do not affect our results significantly). We collect these operators in Table 4.
To calculate low-energy LFV processes, such as muon and tau decays and muon-to-electron conversion in nuclei, we use flavio [40]. For high-energy collider observables, we follow Ref. [67]. As for $0\nu\beta\beta$, the effective Majorana mass $m_{ee}$ is given by $m_{ee}=\left|\sum_{i}m_{i}V^{2}_{ei}\right|$.
To generate predictions for the top FN textures, we retain all trials of our final scan that meet the goodness of fit criterium $\delta_{\mathrm{max}}\leq 2$, storing the associated rotation matrices, $\epsilon$, and, where applicable, $\Lambda_{(W)}$. [^2] We explicitly checked with the top few textures that performing exact fits, as detailed in Appendix A.4, and using these exact fits as input for the phenomenological study did not significantly alter the results. Trials are then ranked by $\delta_{\mathrm{max}}$, and for the top 500, we calculate the suppression for the Wilson coefficients of the operators in Table 4. Note that to remain as model-independent as possible we have excluded operators where the two quark fields have different chiralities, as these could be further suppressed if, for instance, the quarks carried FN charges. However, we explicitly verified that including these operators results in only a $\sim 5\%$ correction. We only included flavor diagonal quark operators for the same reason. After rotating to the mass basis, we derive the predictions for the observables in 3. For each observable we find the average and standard deviation and compare it to data. This allows us to establish a lower bound on $\Lambda$ for Dirac and Majorana scenarios, and to check whether the predictions for type-I seesaw models comply with current bounds.
@@ A.4 Natural Fits
Figure 5: Histograms of the $\eta$ parameter distribution for $1000$ trials that yield exact fits to the measured leptonic masses and mixings parameters. For each neutrino mass mechanism we show the result for the top-performing texture as per Table 1: for Dirac neutrinos (left), $X_{L}=\{6,5,5\}$, $X_{e}=\{3,2,0\}$, $X_{N}=\{-9,-8,-8\}$; for Majorana (middle), $X_{L}=\{2,0,-1\}$, $X_{e}=\{-7,-6,-4\}$; and for type-I seesaw (bottom), $X_{L}=\{6,1,-1\}$, $X_{e}=\{-7,-7,-6\}$, $X_{N}=\{4,0,-3\}$.
@@
For each mass generation mechanism, we focus on the top texture (see Table 1). We use a simplex minimization algorithm to refine the coefficients $c^{\alpha}_{ij}$ to fit the experimental values by minimizing the cost function in Eq. (20). We do this using as starting point 1000 coefficient sets from trials that yielded a promising $\delta_{\mathrm{max}}<2$.
@@ Appendix B More on Phenomenology and Supplemental Plots
Figure 6: Predicted muon collider event rates for CLFV processes in Dirac FN and Majorana FN scenarios with $\Lambda=\Lambda_{\rm exp}\neq\Lambda_{W}$ , using the same assumptions as in Figures 1 and 2, for an integrated luminosity of 10 ab-1. In this figure, the ranking of teach texture is indicated by color, with red textures ranked higher than blue textures. The corresponding results for the Majorana scenario with $\Lambda=\Lambda_{W}$ or type-I seesaw are not shown, as the predicted rates are many orders of magnitude smaller.
@@
This appendix completes the discussion in Section IV, supplementing the results with additional figures.
Low-energy CLFV and bounds on $\Lambda$ For both the Dirac and Majorana cases, the bounds on the FN scale $\Lambda$ are set by current low-energy CLFV bounds. Figure 7 illustrates that the strongest bounds are mostly set by $\mu\to e\gamma$ (and secondly by $\mu\to 3e$), with the strongest bound on $\Lambda$ being remarkably consistent across different textures. The difference between the null and FN textures arises because within FN observables sensitive to left-right couplings are suppressed with respect to the anarchic case.
Future muon colliders – Collider experiments provide a potential platform to test LFV processes predicted by FN models, especially at high energies where muon colliders [50] outshine $e^{+}e^{-}$ colliders [68] due to their higher accessible energies and the resulting enhancement of cross-sections. Our analysis indicates that the Dirac and Majorana scenarios (with $\Lambda=\Lambda_{\text{exp}}$), show the most promising but still very modest potential for detectable signals. Detailed results for these scenarios are illustrated in Fig. 6.
@@
Seesaw phenomenology – The phenomenology of the top 100 FN type-I seesaw models, detailed in Figure 8, aligns broadly with that in the Dirac and Majorana cases (displayed in Figures 1 and 2), with less variability compared to the null texture. Unfortunately, none of the identified textures have a low enough scale to be probed by any currently planned CLFV experiment.
Figure 8: Phenomenology for the top 100 type-I seesaw textures, as in Fig.1. The FN scale $\Lambda$ is fixed by neutrino masses.
Active Neutrino Masses The predicted mass of the lightest neutrino and the sum of neutrino masses is shown in Fig. 9 for the top 100 textures of each mass mechanism. It is interesting to note that many of these natural realistic textures predict an ultralight active neutrino (see e.g. [69, 70]).
Dirac FN only predicts NO neutrinos. While DESI is projected to be able to measure the minimal NO scenario at $3\sigma$ [45], the resolution will not be sufficient to discriminate between Dirac textures. However, a positive determination of IO would strongly disfavor the Dirac scenario.
@@ C.2 Other Sources of Lepton Number Violation
In the type-I seesaw case, we have assumed no additional sources of lepton number violation beyond the Majorana mass of the RH neutrino. If this assumption is relaxed, the Lagrangian in Eq. (9) is modified by an additional contribution to the Weinberg operator:
[…]
The resulting neutrino masses can be derived similarly to those in the type-I seesaw discussed in the main text. We define the Weinberg mass as $m^{2}_{W,ij}=v^{2}c_{ij}^{W}\epsilon^{n_{ij}^{W}}$ and note that the Weinberg operator does not contribute to the mixing between sterile and active eigenstates until $\mathcal{O}\left((v\Lambda^{-1})^{3}\right)$ [71]. The resulting neutrino masses are modified from Eq. (11) to:
[…]
@@
[…]
@@ C.3 Type II Seesaw and other Majorana UV Completions
The Majorana case can arise from a variety of UV completions. For illustration, we consider the example of a FN-augmented type-II seesaw mechanism. The type-II seesaw [35, 36, 37, 38, 39] has the form [72]
@@
$\mathcal{L}_{SS2}\supset-\frac{y_{ij}}{2}\epsilon^{|X_{Li}+X_{Lj}+X_{\Delta}|}L_{i}L_{j}\Delta\\ -c_{\mu}\Lambda\epsilon^{|X_{\Delta}|}\Delta H^{\dagger}H^{\dagger}+\text{h.c.}\\ -(c_{M}\Lambda)^{2}\text{Tr}[\Delta^{\dagger}\Delta].$ (29)
@@
where $\tilde{c}\sim$ $\mathcal{O}(1)$ In the second equality we have assumed all relevant charges to be positive, so that the absolute value can be dropped. As a result, the apparent scale of the Weinberg operator differs from the true FN scale by $\Lambda_{W}=\epsilon^{-2X_{\Delta}}\Lambda$. Consequently, a given effective Weinberg scale can correspond to a much lower physical scale. However, if charges are such that the absolute values in the exponents cannot be dropped, different generations may have distinct FN suppression. This scenario cannot be mapped onto our generic Majorana textures and lies beyond the scope of this work.
@@
$\frac{c^{W}_{2}}{\Lambda_{W}}=$ $\frac{1}{\Lambda}\left[c^{\nu}_{ij}\epsilon^{|X_{L_{i}}+X_{N_{j}}|}\right.$ (31)
$\quad\left.\left(c^{M}_{jm}\epsilon^{|X_{N_{j}}+X_{N_{m}}|}\right)^{-1}c^{\nu}_{mn}\epsilon^{|X_{L_{n}}+X_{N_{Rm}}|}\right]\,.$
@@ Appendix D Auxiliary Material
In the auxiliary material we include several additional documents for reference.
The files

### v2

@@ A.3 Predicting Lepton Violation in FN
For the predictions, we employ the SMEFT framework. We construct effective operators following the power counting dictated by formal $U(1)_{\rm FN}\,$invariance, as in
where $c_{ijkl}$ are ${\cal O}(1)$ coefficients and we have defined
$n_{ijkl}\equiv|X_{\psi_{i}}-X_{\psi_{j}}+X_{\psi_{k}}-X_{\psi_{l}}|.$ (11)
We assume all Wilson coefficients to be $\mathcal{O}(1)$ , and generate them randomly with the same priors used for the effective Yukawa coefficients. For simplicity, we consider only SMEFT operators that contribute at tree-level to our observables (we checked explicitly with wilson [Aebischer:2018bkb] that running effects do not affect our results significantly). We collect these operators in Table 4.
To calculate low-energy LFV processes, such as muon and tau decays and muon-to-electron conversion in nuclei, we use flavio [Straub:2018kue]. For high-energy collider observables, we follow Ref. [Paraskevas:2018mks]. As for $0\nu\beta\beta$, the effective Majorana mass $m_{ee}$ is given by $m_{ee}=\left|\sum_{i}m_{i}V^{2}_{ei}\right|$.
To generate predictions for the top FN textures, we retain all trials of our final scan that meet the goodness of fit criterium $\delta_{\mathrm{max}}\leq 2$, storing the associated rotation matrices, $\epsilon$, and, where applicable, $\Lambda_{(W)}$. [^3] We explicitly checked with the top few textures that performing exact fits, as detailed in Appendix A.4, and using these exact fits as input for the phenomenological study did not significantly alter the results. Trials are then ranked by $\delta_{\mathrm{max}}$, and for the top 500, we calculate the suppression for the Wilson coefficients of the operators in Table 4. Note that to remain as model-independent as possible we have excluded operators where the two quark fields have different chiralities, as these could be further suppressed if, for instance, the quarks carried FN charges. However, we explicitly verified that including these operators results in only a $\sim 5\%$ correction. We only included flavor diagonal quark operators for the same reason. After rotating to the mass basis, we derive the predictions for the observables in 3. For each observable we find the average and standard deviation and compare it to data. This allows us to establish a lower bound on $\Lambda$ for Dirac and Majorana scenarios, and to check whether the predictions for type-I seesaw models comply with current bounds. Note that our bounds on $\Lambda$ are derived purely from processes involving SM states. Considering the phenomenology of the flavon itself may yield additional, but model-dependent, bounds [Heikinheimo:2018luc].
@@ A.4 Natural Fits
Figure 5: Histograms of the $\eta$ parameter distribution for 10001000 trials that yield exact fits to the measured leptonic masses and mixings parameters. For each neutrino mass mechanism we show the result for the top-performing textures: for Dirac neutrinos (left), $X_{L}=\{6,5,5\}$, $X_{e}=\{-3,-2,0\}$, $X_{N}=\{9,8,8\}$; for Majorana (middle), $X_{L}=\{2,0,-1\}$, $X_{e}=\{7,6,4\}$; and for type-I seesaw (bottom), $X_{L}=\{6,1,-1\}$, $X_{e}=\{7,7,6\}$, $X_{N}=\{3,0,-4\}$.
@@
For each mass generation mechanism, we focus on the top texture as ranked in the main text. We use a simplex minimization algorithm to refine the coefficients $c^{\alpha}_{ij}$ to fit the experimental values by minimizing the cost function in Eq. (9). We do this using as starting point 1000 coefficient sets from trials that yielded a promising $\delta_{\mathrm{max}}<2$.
@@ Appendix B More on Phenomenology and Supplemental Plots
Figure 7: Predicted muon collider event rates for CLFV processes in Dirac FN and Majorana FN scenarios with $\Lambda=\Lambda_{\rm exp}\neq\Lambda_{W}$ chosen to saturate current experimental bounds at $\Lambda\sim 10^{6}$ GeV. assuming an integrated luminosity of 10 ab-1. In this figure, the ranking of each texture is indicated by color, with red textures ranked higher than blue textures. The corresponding results for the Majorana scenario with $\Lambda=\Lambda_{W}$ or type-I seesaw are not shown, as the predicted rates are many orders of magnitude smaller.
This appendix completes the discussion in the main text, supplementing the results with additional figures.
Low-energy CLFV and bounds on $\Lambda$ For both the Dirac and Majorana cases, the bounds on the FN scale $\Lambda$ are set by current low-energy CLFV bounds. Fig. 6 illustrates that the strongest bounds are mostly set by $\mu\to e\gamma$ (and secondly by $\mu\to 3e$), with the strongest bound on $\Lambda$ being remarkably consistent across different textures. The difference between the null and FN textures arises because FN observables sensitive to left-right couplings are suppressed with respect to the anarchic case.
Future muon colliders – Collider experiments provide a potential platform to test LFV processes predicted by FN models, especially at high energies where muon colliders [Accettura:2023ked] outshine $e^{+}e^{-}$ colliders [Shiltsev:2019rfl] due to their higher accessible energies and the resulting enhancement of cross-sections. Our analysis indicates that the Dirac and Majorana scenarios (with $\Lambda=\Lambda_{\text{exp}}$), show the most promising but still very modest potential for detectable signals. Detailed results for these scenarios are illustrated in Fig. 7.
@@
Seesaw phenomenology – The phenomenology of the top 100 FN type-I seesaw models, detailed in Fig. 8, aligns broadly with that in the Dirac and Majorana cases (displayed in the main text), with less variability compared to the null texture. Unfortunately, none of the identified textures have a low enough scale to be probed by any currently planned CLFV experiment.
Figure 8: Average predicted CLFV decay rates for the 100 most realistic natural type-I seesaw FN textures (gray lines), relative to each observable’s current constraint. The FN scale $\Lambda$ is fixed by neutrino masses. . Green shading indicates the reach of proposed future low-energy CLFV experiments, and the flavor-anarchic null texture is shown as a red line for comparison.
Active Neutrino Masses The predicted mass of the lightest neutrino and the sum of neutrino masses is shown in Fig. 9 for the top 100 textures of each mass mechanism. It is interesting to note that many of these natural realistic textures predict an ultralight active neutrino (see e.g. [Chacko:2016hvu, Alonso-Alvarez:2023bat]).
Dirac FN only predicts normal-ordered (NO) neutrinos. While DESI is projected to be able to measure the minimal NO scenario at $3\sigma$ [DESI:2016fyo], the resolution will not be sufficient to discriminate between Dirac textures. However, a positive determination of inverted-ordering (IO) would strongly disfavor the Dirac scenario.
@@ C.2 Other Sources of Lepton Number Violation
In the type-I seesaw case, we have assumed no additional sources of lepton number violation beyond the Majorana mass of the RH neutrino. If this assumption is relaxed, the Lagrangian is modified by an additional contribution to the Weinberg operator:
$\mathcal{L}\supset\mathcal{L}_{\rm SS}-\frac{c_{ij}^{W}\epsilon^{n^{W}_{ij}}}{\Lambda_{\cancel{L}}}(L_{i}H)(L_{j}H)\,$, (16)
where $\Lambda_{\cancel{L}}$ indicates the scale of the additional lepton number violation.
The resulting neutrino masses can be derived similarly to those in the type-I seesaw discussed in the main text. We define the Weinberg mass as $m^{2}_{W,ij}=v^{2}c_{ij}^{W}\epsilon^{n_{ij}^{W}}$ and note that the Weinberg operator does not contribute to the mixing between sterile and active eigenstates until $\mathcal{O}\left((v\Lambda^{-1})^{3}\right)$ [Dedes:2006ni]. The resulting neutrino masses are:
$m_{\nu}\approx\frac{m^{2}_{W}}{\Lambda_{\cancel{L}}}-m_{D}M_{N}^{-1}m_{D}^{T}\,$, (17)
@@
We performed our analysis in the case where $\Lambda_{\cancel{L}}=\Lambda=M$, and found slightly different top textures. However, the overall conclusions remain unchanged: the scale of flavor violation for these models exceeds projected experimental sensitivities. The CLFV phenomenology for the top 100 textures is shown in Fig. 10.
Figure 10: Phenomenology for the top 100 type-I seesaw textures if the FN sector is lepton number violating (with $\Lambda_{\cancel{L}}=\Lambda=M$), as in Fig. 8. The FN scale $\Lambda$ is fixed by neutrino masses. The axes are fixed to be the same as in Fig. 8 for ease of comparison.
@@ C.3 Type II Seesaw and other Majorana UV Completions
The Majorana case can arise from a variety of UV completions. For illustration, we consider the example of a FN-augmented type-II seesaw mechanism. The type-II seesaw [Magg:1980ut, Schechter:1980gr, Lazarides:1980nt, Mohapatra:1980yp, Wetterich:1981bx] has the form [Ma:1998dx]
@@
$\mathcal{L}_{SS2}\supset-\frac{y_{ij}}{2}\epsilon^{|X_{Li}+X_{Lj}+X_{\Delta}|}L_{i}L_{j}\Delta\\ -c_{\mu}\Lambda\epsilon^{|X_{\Delta}|}\Delta H^{\dagger}H^{\dagger}\\ +\text{h.c.}-(c_{M}\Lambda)^{2}\text{Tr}[\Delta^{\dagger}\Delta].$ (20)
@@
where $\tilde{c}\sim$ $\mathcal{O}(1)$ . In the second equality we have assumed all relevant charges to be positive, so that the absolute value can be dropped. As a result, the apparent scale of the Weinberg operator differs from the true FN scale by $\Lambda_{W}=\epsilon^{-2X_{\Delta}}\Lambda$. Consequently, a given effective Weinberg scale can correspond to a much lower physical scale. However, if charges are such that the absolute values in the exponents cannot be dropped, different generations may have distinct FN suppression. This scenario cannot be mapped onto our generic Majorana textures and lies beyond the scope of this work.
@@
$\frac{c^{W}_{2}}{\Lambda_{W}}=\frac{1}{\Lambda}\left[c^{\nu}_{ij}\epsilon^{|X_{L_{i}}+X_{N_{j}}|}\right.\\ \left.\left(c^{M}_{jm}\epsilon^{|X_{N_{j}}+X_{N_{m}}|}\right)^{-1}c^{\nu}_{mn}\epsilon^{|X_{L_{n}}+X_{N_{Rm}}|}\right]\,.$ (22)
@@ Appendix D Auxiliary Material
In the auxiliary material we include several additional documents for reference. The files